OpenStudy (anonymous):

If x varies inversely with y and x = 2 when y = 48, find x when y = 24. A)2 B)4 C)1 D)12

jimthompson5910 (jim_thompson5910):

x varies inversely with y means that x = k/y where k is some fixed number

jimthompson5910 (jim_thompson5910):

it tells us that when x = 2, y is y = 48, so this means x = k/y 2 = k/48 .. plug in x = 2 and y = 48 now solve for k and tell me what you get

OpenStudy (anonymous):

k=96

jimthompson5910 (jim_thompson5910):

so your equation is now y = 96/x

jimthompson5910 (jim_thompson5910):

when x = 24, y = ???

OpenStudy (anonymous):

y=4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

oh wait, sorry x = 96/y is your equation x = 96/24 x = 4 the answer is still 4 though

OpenStudy (anonymous):

ah! sweet! thank you so much for helping me!!! :)

jimthompson5910 (jim_thompson5910):

you're welcome